OpenStudy (anonymous):

I have algebra and cant figure out this problem 72+4-14d=36

OpenStudy (♪chibiterasu):

Basically they are asking you to isolate d. We do this by canceling out certain values from both sides of the equation in order to get x alone. First, we subtract 72 from both sides in order to remove it from the left side: 72 - 72 + 4 - 14d = 36 - 72 4 - 14d = -36 Do you know what we do next?

OpenStudy (anonymous):

My teacher said to divide but i dont know were to divide

OpenStudy (♪chibiterasu):

Well we don't do that yet, we have to get rid of the 4. How?

OpenStudy (anonymous):

Add 4 to both side 4 - 4 - 14d = -36 - 4

OpenStudy (♪chibiterasu):

Well subtract, but you still wrote the equation right anyway :P 14d = -40 Now we get to dividing. What do we divide now?

OpenStudy (anonymous):

divide 14d and 14 they will cancel each other divide 14d and -40 d = -0.35

OpenStudy (♪chibiterasu):

Mm, not quite. You divide 14 yes, but the end quotient isn't -0.35 . . .

OpenStudy (♪chibiterasu):

-40/14 isn't -0.35 :c

OpenStudy (anonymous):

would it be a fraction

OpenStudy (♪chibiterasu):

No, it's still a number.

OpenStudy (anonymous):

The easy way to look at it for any equation, no matter how high your math level is: What you do on one side of an equation (equals sign) you must do on the other side of the equation. Therefore, in this case, since you are trying to get 'd' all by itself, first make sure that the term containing 'd' is isolated. Do so by adding or subtracting all of the values that are on one side (and do the same to the other side) leaving the '14d' term on its own on one side, and a purely numerical value on the other side. Once you have done this, divide BOTH sides by 14, leaving '1d' which equals 'd' by itself on one side, and the other side will be just a number.

OpenStudy (♪chibiterasu):

We're already doing that. but thank you for your description! :3

OpenStudy (♪chibiterasu):

We're on the last step, try popping out a calculator and solve -40/14 o:

OpenStudy (anonymous):

is it a decimal or a whole number

OpenStudy (♪chibiterasu):

Decimal, but it's infinite so it needs to be rounded.

OpenStudy (♪chibiterasu):

It should be rounded to the next two digits, as in the nearest 100th

OpenStudy (anonymous):

it says it is 2.85... so it would be rounded to 3?

OpenStudy (♪chibiterasu):

well yes, the answer is 2.85 but it's not positive :P

OpenStudy (anonymous):

-3

OpenStudy (♪chibiterasu):

You don't round it to the whole number, you round it to the nearest 100th so it would be -2.85 :3

OpenStudy (anonymous):

thank u. u helped me alot

OpenStudy (♪chibiterasu):

Your welcome, any time :3